CAT Chat Minutes

February 7, 2005

 

Announcements and Follow-up
Bike Storage– E. Chang is looking into re-arranging cabinets to accommodate bike storage.   He is on vacation today.

Mail to Sectors– Steve Downey said that it really is a budget issue.   Kevin Randall offered to check if additional mail drops could be made on the opposite end of the ring.

User reported that Users from ORNL are having trouble calling APS.  

Kevin Randall will see what is happening with the phones.

Machine Stats
Statistics for Run1-2005

Delivered beam:   165.11 hrs.
Percentage of scheduled time:   97.65%
Downtime during period: 3.97 hrs.
Mean time between faults: 165.11 hrs.

Faults:   Listing of Statistics for Run 1-2005

Fill
Reason
Downtime
#   1   01/31 08:00   To   02/01 03:59    19.99 S24-ID Slow Valve   [XFD]   3.97 2 hrs., 40min=XFD. 1hr. 17 mins.=PSG S24 canted dipole.
#   2   02/01 07:58   To   02/07 10:04   146.11 Fill Ongoing 0

Beamline Data
Beamtime downtime summary by APS groups, 01/28/2005 to 02/4/2005

DIN

Date

Beamline

Status

Duration

Group

Description

Resolution

493

01/31/05

19-ID

Closed

3 hr 48 min

ASD-SI

Chain B Love controller for Station C White Beam Stop is not working properly.

Love controller replaced 1-for-1. New Love controller was validated to confirm proper operation.

494

01/31/05

16-ID

Closed

20 min

ASD-CTL

P. Chow reported a timer counting down on his ID which opens the ID all the way. This function should not run during User Operations with the shutter open. M. Smith is investigating. T. Gog (9-ID) reported the same problem, his ID auto opened to 30 mm twice today.

Gap timer function reset for all IDs by ASD-CTL. Proper operation of the controls software has been restored.

496

01/31/05

24-ID

Closed

20 hr 0 min

XFD-XFE

SV does not respond to open command

Storage Ring access found Slow Valve control switch set to "local" instead of "remote". Setting changed to remote and control of the SV was restored.

497

02/01/05

1-BM

Closed

5 hr 4 min

ASD-VAC

Serious Fault #160 (P7C_MS1_NOSWITCH) occurred. Joe Gagliano submitted Work Request # 16942 to rebuild the shutter.

ASD-VAC rebuilt MS1. The problem was with the seals. New seals installed and shutter put back together. After repair, ASD-SI performed mini-validation on component. Beamline restored to proper operations.

Current schedule: http://www.aps.anl.gov/Facility/Schedule/

Safety
Jim Lang, ASD Safety, recapped events of the electrical stand-down and said by the evening of January 28 th, APS division and on most APS beamlines had been able to restart electrical work.